# DP Technology *

### v1

@@ 6 Applications
Document Understanding. Uni-Parser significantly enhances scientific literature workflows by supporting automatic document summarization, document-level question answering [cai2024uni], paper-to-poster [pang2025paper2poster] and paper-to-PPT [dataflow-agent2025] generation, as well as intelligent retrieval and deep research [openai2025deepresearch]. Collectively, these capabilities streamline both knowledge consumption and dissemination for researchers across diverse disciplines.
Structured Data Extraction. By converting unstructured documents into structured representations, Uni-Parser supports the large-scale construction of domain-specific databases, such as paper citation database, scholar database, molecular libraries, reaction repositories, bioactivity database [yan2025biominer], experimental characterization database [wang2025uniem3muniversalelectronmicrograph, xu2025toward], and comprehensive entity knowledge bases. Such resources are critical for accelerating data-driven scientific discovery.
Patent Retrieval and Protection. The framework further facilitates patent retrieval and prior-art verification, providing robust support for innovation discovery and intellectual property protection [shi2024intelligent, zhuang2025doc2sar]. This enables more efficient navigation of complex patent landscapes in scientific and industrial settings.
Foundation Model Training. Finally, Uni-Parser can serve as a powerful engine for large-scale training data generation, supplying high-quality structured inputs that reduce the cost and effort of manual curation. This capability is particularly valuable for advancing foundation models in scientific domains, where large, reliable datasets are indispensable [li2024scilitllm, liao2025innovator, dataflow2025].
@@ 8 Future Work
Uni-Parser-Tools for Easy Access. We will release a fully packaged, open-source toolkit, Uni-Parser-Tools, which enables remote access to Uni-Parser without requiring local computational resources. This toolkit will include example pipelines for downstream applications, allowing users to quickly leverage Uni-Parser to construct structured scientific databases and unlock generative AI applications for scientific discovery.
Benchmark Construction. Existing benchmarks, such as OmniDocBench and our Uni-Parser benchmark, are constrained by inconsistencies in layout recognition and the heterogeneous output formats of generalized OCR modules. Moreover, they struggle to adequately represent complex layouts and cross-page content. We aim to investigate more robust approaches to benchmark construction, particularly task-driven benchmarks tailored to downstream applications, including document understanding and structured data extraction. Such efforts will enable more fair, comprehensive, and informative evaluations of parsing performance.
@@ Appendix A Authorship and Acknowledgments
Program Leads [^1] Contributors listed in alphabetized order.
@@ Appendix B Examples and Comparisons
B.2 Uni-Parser Qualitative Examples
Examples here.

### v2

@@ 6 Applications
Document Understanding. Uni-Parser significantly enhances scientific literature workflows by supporting automatic document summarization, document-level question answering [3], paper-to-poster [30] and paper-to-PPT [40] generation, as well as intelligent retrieval and deep research [27]. Collectively, these capabilities streamline both knowledge consumption and dissemination for researchers across diverse disciplines.
Structured Data Extraction. By converting unstructured documents into structured representations, Uni-Parser supports the large-scale construction of domain-specific databases, such as paper citation database, scholar database, molecular libraries, reaction repositories, bioactivity database [52], experimental characterization database [47, 50, 44], and comprehensive entity knowledge bases. Such resources are critical for accelerating data-driven scientific discovery.
Patent Retrieval and Protection. The framework further facilitates patent retrieval and prior-art verification, providing robust support for innovation discovery and intellectual property protection [37, 57]. This enables more efficient navigation of complex patent landscapes in scientific and industrial settings.
Foundation Model Training. Finally, Uni-Parser can serve as a powerful engine for large-scale training data generation, supplying high-quality structured inputs that reduce the cost and effort of manual curation. This capability is particularly valuable for advancing foundation models in scientific domains, where large, reliable datasets are indispensable [14, 16, 39].
@@ 8 Future Work
Uni-Parser-Tools for Easy Access. We will release an open-source toolkit, Uni-Parser-Tools, providing remote access to Uni-Parser without requiring local compute. The toolkit includes example pipelines for downstream tasks, enabling rapid construction of structured scientific databases and supporting generative AI applications for scientific discovery.
Benchmark Construction. Existing benchmarks (e.g., OmniDocBench and the Uni-Parser benchmark) are limited by inconsistent layout annotations and heterogeneous OCR outputs, and fail to adequately capture complex layouts and cross-page content. We plan to explore more robust, task-driven benchmarks tailored to downstream applications such as document understanding and structured data extraction, enabling fairer and more informative evaluations.
@@ Appendix A Authorship and Acknowledgments
Team Management
@@ Appendix B Examples and Comparisons
B.2 Qualitative Examples of Cross-Page Layout Grouping
B.3 Layout Detection Comparison
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
B.4 Qualitative Examples of Reading Order
Original Document
Uni-Parser
Original Document
Uni-Parser
Original Document
Uni-Parser
Original Document
Uni-Parser
B.5 Uni-Parser Case Study
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5
Original Document
Uni-Parser
$P$-StructureV3
MinerU2.5